# Generalizing CLIP to Unseen Domain via Text-Guided Diverse Novel Feature Synthesis

### v1

@@ 4.2 Evaluation
Table 3: Top 1 accuracy % on DomainNet (ViT-B/16).
@@
Main results. First, we benchmark various SOTA algorithms and our LDFS on PACS, OfficeHome, DomainNet, and NICO++. The results are reported in Table 1, 9, 3, and 4. The main observations from existing algorithms include: 1) We notice linear probing, i.e.,CLIP (LP), can hurt the generalization ability of the pre-trained CLIP model, as demonstrated by its subpar performance compared to zero-shot CLIP, i.e.,CLIP (ZS) across all four datasets. 2) Existing CLIP-based finetuning methods, along with text-guided feature synthesis methods such as DPL, LADS, and PODA, are not always effective for target domain adaptation. For instance, both DPL and LADS degrade the average performance of zero-shot CLIP in PACS and NICO++, as shown in Table 1 and 4.
Table 4: Top 1 accuracy % on NICO++ (ResNet50).
@@
Table 5: Synthetic feature quality (%)
on PACS
@@
Table 6: Comparison (%) with generative models
Method Art Cartoon Sketch Avg.
CLIP (ZS) 97.23 99.10 87.81 94.77
CoOp + CLIPStyler 96.33 98.19 89.06 94.53
CoOp + CLIPStyler† 95.38 97.52 89.15 94.02
CoOp + LDFS 97.84 99.19 93.62 96.88
Further, adding our LDFS to linear probing or current prompt learning techniques like CoOp, CoCoOp, MaPle, and PromptSRC results in significant performance improvements, as highlighted by blue numbers in Table 1 to 4. Specifically, our LDFS combined with PromptSRC outperforms all baselines on average accuracy across all datasets while also achieving the best accuracy on most target domains. Also, it improves PromptSRC alone by 2.02%, 1.78%, 2.36%, and 1.71% respectively on four datasets. Additionally, LDFS significantly improves the efficiency of existing prompt learning techniques for certain domains. For example, CoOp + LDFS achieves a 3.11% improvement over CoOp in the Art domain of OfficeHome and a 4.17% improvement over CoOp in the Quickdraw domain of DomainNet. Similarly, MaPle + LDFS improves over MaPle by 4.58% on the Grass domain of NICO++. In the linear probing setting, LDFS (LP+LDFS) outperforms LADS and PODA on average accuracy on NICO++. Notably, since PODA is specifically designed for CNN architectures, no PODA results are available beyond Table 4.
@@
Figure 5: Style transfer photo images into other domains via CLIPStyler [17]. Unexpected artifacts (red box) and the model’s inability to synthesize target objects (blue box) hinder its efficacy for downstream adaptation.
[…]
Analysis of the synthetic features quality. We assess the quality of synthetic features in our LDFS against the SOTA text-guided feature synthesis method, LADS [5]. On the PACS dataset, given photo domain images, we synthesize novel features for other domains using both methods. Since neither method directly generates images, we employ nearest neighbor images from the respective target domains to verify synthetic feature quality. In Fig. 4, we display 8 randomly chosen samples from the photo domain and the corresponding synthetic target domain images. It shows LADS and LDFS successfully generate target domain features. Nevertheless, LADS often produces repetitive or similar features for the same category, as highlighted in the red boxes of Fig. 4 (a). Meanwhile, LADS struggles to preserve class information, especially for those intricate samples during adaptation, as shown in green boxes of Fig. 4 (a). In contrast to LADS, our LDFS can synthesize more diverse target domain features with superior quality while maintaining class consistency, as shown in Fig. 4 (b). Furthermore, Table 12 shows that LDFS achieves the highest domain alignment (DA), class consistency (CC), and diversity scores (DS), which further demonstrates the efficacy of our strategy. More quantitative and qualitative results, including details on these scores, are provided in Appendix 1.4 and 2.
4.3 Ablation Studies
@@
In this subsection, we focus on ablating the major components of our proposed framework. We provide additional ablations, including analyses of different text generation strategies, domain text descriptions, number of training images, and more results of synthetic feature quality, in Appendix. Also, when we refer to LDFS during ablation study, we are specifically referring to the LDFS + CoOp.
@@
Analysis of stochastic text feature augmentation. We carry out experiments to investigate the effectiveness of our introduced stochastic text feature augmentation approach. Specifically, we analyze the effect of the noise level, denoted as $\gamma$, on the modality gap and subsequent performance. As depicted in the left panel of Fig. 6, there’s a consistent decline in the modality gap as the noise level rises, demonstrating our method’s ability to bridge this gap. Correspondingly, the right panel of Fig. 6 illustrates that by closing the modality gap via our technique, LDFS performance is enhanced in target domain adaptation. For most noise levels, stochastic text feature augmentation can enhance LDFS performance compared to when there’s no augmentation (i.e.,noise level=0). The peak performance boost is observed at a noise level of 0.1 across all three domains. Additionally, referring to Table 4.3, it’s evident that our method (i.e.,LDFS), when using tuned $\gamma$, boosts LDFS performance without augmentation (i.e.,LDFS w/o TA) by respective margins of 2.99%, 1.51%, and 0.31% across PACS, OfficeHome, and NICO++ benchmarks. Compared to increasing the gap (LDFS w/TA+) and adding noise to image features (LDFS w/ IA), our text augmentation method (LDFS) performs the best.
@@
Table 7: Ablation studies on different datasets (%).
@@
Figure 8: Sensitivity analysis of the weighting hyperparameters $\alpha$ and $\beta$. The red and blue dash lines represent the performance of the model without using LDFS.
@@
Ablation on the loss. In Table 4.3, we evaluate the effect of $\mathcal{L}_{ld}$, $\mathcal{L}_{pair}$, and $\mathcal{L}_{cc}$ used on our LDFS. LDFS w/ global denotes our baseline modified by replacing $\mathcal{L}_{ld}$ with the conventional global adaptation loss [5]. We can see that excluding the local adaptation losses $\mathcal{L}_{ld}$ or $\mathcal{L}_{pair}$ leads to performance degradation across all datasets. Also, removing $\mathcal{L}_{cc}$ results in significant degradation, indicating the importance of preserving category information during feature synthesis. Moreover, LDFS, when fully equipped with all components, outperforms all baselines across all datasets, showcasing the importance of all proposed components. From an alternative perspective on synthetic feature quality, we study the ablation results of translating features from the photo domain into the art domain, as illustrated in Fig. 7. Red boxes highlight the issue of homogeneous feature synthesis, the blue box indicates models’ failures in domain translation, and the green boxes show models struggle to preserve class information. $\mathcal{L}_{da}$ indicates baseline with global adaptation loss. We observe that replacing $\mathcal{L}_{da}$ with our local adaptation loss $\mathcal{L}_{ld}$ can address the diversity issue, yet result in the synthesis of incorrect class or domain features. Combining all components together leads to the best quality of synthetic features.
Analysis of Instance-conditional Loss. We also compared our loss variant using text description produced by GPT-3 in Table 5 in Appendix. It shows LDFS†(w/ GPT-3) outperforms LADS (LDFS w/ global text) but underperforms LDFS, showing captioning model can produce more aligned text description than using the GPT-3 strategy. We also tried averaging random descriptions following WaffleCLIP [30] but found no clear improvement over single description. Overall, we observe that better semantic alignment in image-text leads to more performance boosts, whereas simply adding or averaging stochastic directions does not result in as much improvement.
@@
The Effect of Text Description. To explore if our method relies on users’ verbalization skills, we ablate different text descriptions in Table 4 of the Appendix. It shows consistent efficacy despite variations in description format. Moreover, Table 2 in the Appendix displays straightforward, uniform text descriptions used across all our datasets in our paper. The simplicity of these descriptions demonstrates that our method is easily applicable and user-friendly.

### v2

@@ IV-B Evaluation
Main results. First, we benchmark various SOTA algorithms and our LDFS on PACS, OfficeHome, DomainNet, and NICO++. The results are reported in Table II, III, IV, and V. The main observations from existing algorithms include: 1) We notice linear probing, i.e. CLIP (LP), can hurt the generalization ability of the pre-trained CLIP model, as demonstrated by its subpar performance compared to zero-shot CLIP, i.e. CLIP (ZS) across all four datasets. 2) Existing CLIP-based finetuning methods, along with text-guided feature synthesis methods such as DPL, LADS, and PODA, are not always effective for target domain adaptation. For instance, both DPL and LADS degrade the average performance of zero-shot CLIP in PACS and NICO++, as shown in Table II and V.
TABLE IV: Top 1 accuracy % on DomainNet (ViT-B/16).
@@
TABLE V: Top 1 accuracy % on NICO++ (ResNet50).
@@
TABLE VI: Comparison (%) with the generative models.
Method Art Cartoon Sketch Avg.
CLIP (ZS) [29] 97.23 99.10 87.81 94.77
CoOp + CLIPStyler [17] 96.33 98.19 89.06 94.53
CoOp + CLIPStyler†[17] 95.38 97.52 89.15 94.02
CoOp + LDFS 97.84 99.19 93.62 96.88
TABLE VII: Synthetic feature quality (%) on PACS .
@@
Further, adding our LDFS to linear probing or current prompt learning techniques like CoOp, CoCoOp, MaPle, and PromptSRC results in significant performance improvements, as highlighted by blue numbers in Table II to V. Specifically, our LDFS combined with PromptSRC outperforms all baselines on average accuracy across all datasets while also achieving the best accuracy on most target domains. Also, it improves PromptSRC alone by 2.02%, 1.78%, 2.36%, and 1.71% respectively on four datasets. Additionally, LDFS significantly improves the efficiency of existing prompt learning techniques for certain domains. For example, CoOp + LDFS achieves a 3.11% improvement over CoOp in the Art domain of OfficeHome and a 4.17% improvement over CoOp in the Quickdraw domain of DomainNet. Similarly, MaPle + LDFS improves over MaPle by 4.58% on the Grass domain of NICO++. In the linear probing setting, LDFS (LP+LDFS) outperforms LADS and PODA on average accuracy on NICO++. Notably, since PODA is specifically designed for CNN architectures, no PODA results are available beyond Table V.
Refer to caption Refer to caption
(a) Synthesized feature nearest neighbors using LADS [5]. (b) Synthesized feature nearest neighbors using LDFS (ours).
Figure 5: Style transfer photo images into art, cartoon, and sketch domains via CLIPStyler [17]. Unexpected artifacts (red box) and the model’s inability to synthesize target objects (blue box) hinder its efficacy for downstream adaptation.
Refer to caption Refer to caption
[…]
Analysis of the synthetic features quality on PACS. We assess the quality of synthetic features in our LDFS against the SOTA text-guided feature synthesis method, LADS [5]. On the PACS dataset, given photo domain images, we synthesize novel features for other domains using both methods. Since neither method directly generates images, we employ nearest neighbor images from the respective target domains to verify synthetic feature quality. In Fig. 4, we display 8 randomly chosen samples from the photo domain and the corresponding synthetic target domain images. It shows LADS and LDFS successfully generate target domain features. Nevertheless, LADS often produces repetitive or similar features for the same category, as highlighted in the red boxes of Fig. 4 (a). Meanwhile, LADS struggles to preserve class information, especially for those intricate samples during adaptation, as shown in green boxes of Fig. 4 (a). In contrast to LADS, our LDFS can synthesize more diverse target domain features with superior quality while maintaining class consistency, as shown in Fig. 4 (b). Furthermore, to quantitatively evaluate the quality of the synthesized features, we follow [5] in using the domain alignment (DA) score to evaluate the model’s ability to correctly translate data to the target domain. We also use the class consistency (CC) score to evaluate how well the model preserves class information during synthesis. Further, we propose a feature diversity (DS) score to evaluate the diversity of the synthesized features. Specifically, the DA score is obtained by calculating the percentage of nearest neighbors belonging to the desired target domain. The CC score is obtained by calculating the percentage of synthetic features that belong to the original class. The DS score is obtained by one minus the percentage of repeated nearest neighbors within the same class. Table VI shows that LDFS achieves the highest domain alignment (DA), class consistency (CC), and diversity scores (DS), which further demonstrates the efficacy of our strategy.
@@
Figure 7: Comparison of nearest neighbor images of synthesized target domain features between LADS and our LDFS on the NICO++ dataset: outdoor → autumn, dim, grass, rock, and water (zoom in for better visualisation). The first row displays randomly sampled images from the source domain (outdoor). Subsequent rows showcase the nearest neighbors of synthetic features from both LADS and LDFS. Red boxes highlight issues with homogeneous synthetic features lacking diversity for both methods. Blue boxes indicate where both methods fail to generate target domain features. Green boxes show where both lose class information during adaptation, although LDFS preserves more class information than LADS.
Analysis of the synthetic features quality on NICO++. Additionally, we compare the synthetic feature quality of our LDFS against LADS [5] on NICO++ dataset [41]. Given outdoor domain images, we synthesize novel features for the autumn, dim, grass, rock, and water domains using both LDFS and LADS. In Fig. 7, we display 8 randomly chosen samples from the outdoor domain and corresponding target domains. Both LDFS and LADS are effective for synthesizing target domain features in most cases. However, LADS often produces repetitive features lacking diversity (red boxes in Fig. 7 (a)) or wrong class features (green boxes in Fig. 7 (b)). In contrast, our LDFS synthesizes more diverse, high-quality target domain features while maintaining class consistency. However, we observe both LDFS and LADS can sometimes generate wrong domain features (blue boxes in Fig. 7). This rarely occurred on the simpler PACS dataset, indicating the NICO++ domains are more complex and challenging. Moreover, Table VIII shows that our LDFS achieves the highest class consistency (CC) and diversity scores (DS), but a slightly lower domain alignment score (DA) compared to LADS [5]. Furthermore, it can be seen that the generative model CLIPStyler [17] achieves a higher diversity score than LADS, but much lower domain alignment and class consistency scores.
TABLE VIII: Synthetic feature quality (%) on NICO++ .
LADS CLIPStyler LDFS
DA score 96.18 74.27 94.33
CC score 82.96 68.69 88.54
DS score 76.86 85.81 92.45
TABLE IX: Ablation on the effect of different text descriptions over all domains on NICO++ [41].
Base Text Description CoOp [45] CoOp+LDFS
“a photo of a {class name} with {domain name} background.” 81.95 83.23 (+1.28)
“a {domain name} style photo of a {class name}.” 81.95 82.93 (+0.98)
“a {domain name} background photo of a {class name}.” 81.95 83.20 (+1.25)
TABLE X: Base text descriptions used in different datasets.
PACS OfficeHome DomainNet NICO++
Source Prompts “a real photo of a {}.” “a real photo of a {}.” “a realistic photo of a {}.” “a photo of a {} with outdoor background.”
Target Prompts “a art photo of a {}.” “a clipart photo of a {}.” “a sketch photo of a {}.” “a photo of a {} with autumn background.”
“a sketch photo of a {}.” “a stock photo of a {}.” “a painting photo of a {}.” “a photo of a {} with dim background.”
“a cartoon photo of a {}.” “a sketch photo of a {}.” “a clipart photo of a {}.” “a photo of a {} with grass background.”
“a painting photo of a {}.” “a art photo of a {}.” “a infograph photo of a {}.” “a photo of a {} with rock background.”
- - “a quickdraw photo of a {}.” “a photo of a {} with water background.”
IV-C Ablation Studies
@@
In this subsection, we focus on ablating the major components of our proposed framework. We provide additional ablations, including analyses of different text generation strategies, domain text descriptions, number of training images, synthetic feature quality, etc. Also, when we refer to LDFS during ablation study, we are specifically referring to the LDFS + CoOp.
@@
Analysis of stochastic text feature augmentation. We carry out experiments to investigate the effectiveness of our introduced stochastic text feature augmentation approach. Specifically, we analyze the effect of the noise level, denoted as $\gamma$, on the modality gap and subsequent performance. As depicted in the left panel of Fig. 6, there’s a consistent decline in the modality gap as the noise level rises, demonstrating our method’s ability to bridge this gap. Correspondingly, the right panel of Fig. 6 illustrates that by closing the modality gap via our technique, LDFS performance is enhanced in target domain adaptation. For most noise levels, stochastic text feature augmentation can enhance LDFS performance compared to when there’s no augmentation (i.e., noise level=0). The peak performance boost is observed at a noise level of 0.1 across all three domains. Additionally, referring to Table XI, it’s evident that our method (i.e., LDFS), when using tuned $\gamma$, boosts LDFS performance without augmentation (i.e., LDFS w/o TA) by respective margins of 2.99%, 1.51%, and 0.31% across PACS, OfficeHome, and NICO++ benchmarks. Compared to increasing the gap (LDFS w/TA+) and adding noise to image features (LDFS w/ IA), our text augmentation method (LDFS) performs the best.
@@
TABLE XI: Ablation studies on different datasets (%).
@@
Refer to caption Refer to caption
XX(a) $\alpha$ XX(b) $\beta$
@@
Figure 9: Sensitivity analysis of the weighting hyperparameters $\alpha$ and $\beta$. The red and blue dash lines represent the performance of the model without using LDFS on PACS and OfficeHome.
@@
Ablation on the loss. In Table XI, we evaluate the effect of $\mathcal{L}_{ld}$, $\mathcal{L}_{pair}$, and $\mathcal{L}_{cc}$ used on our LDFS. LDFS w/ global denotes our baseline modified by replacing $\mathcal{L}_{ld}$ with the conventional global adaptation loss [5]. We can see that excluding the local adaptation losses $\mathcal{L}_{ld}$ or $\mathcal{L}_{pair}$ leads to performance degradation across all datasets. Also, removing $\mathcal{L}_{cc}$ results in significant degradation, indicating the importance of preserving category information during feature synthesis. Moreover, LDFS, when fully equipped with all components, outperforms all baselines across all datasets, showcasing the importance of all proposed components. From an alternative perspective on synthetic feature quality, we study the ablation results of translating features from the photo domain into the art domain, as illustrated in Fig. 8. Red boxes highlight the issue of homogeneous feature synthesis, the blue box indicates models’ failures in domain translation, and the green boxes show models struggle to preserve class information. $\mathcal{L}_{da}$ indicates baseline with global adaptation loss. We observe that replacing $\mathcal{L}_{da}$ with our local adaptation loss $\mathcal{L}_{ld}$ can address the diversity issue, yet result in the synthesis of incorrect class or domain features. Combining all components together leads to the best quality of synthetic features. Similar observations can also be found on NICO++ dataset, as shown in Fig. 10.
Figure 10: Nereast neighbor images of our LDFS when ablating different components on NICO++. $\mathcal{L}_{da}$ denotes global direction adaptation loss of LADS. Red boxes show the issue of homogenous synthetic features; blue boxes show the model fails to synthetic target domain features; and green boxes show the model fails to preserve class information during adaptation.
TABLE XII: Ablation on Different descriptions(%).
Method PACS OfficeHome NICO++
LDFS 96.51 82.96 83.23
LDFS_en 95.97 82.40 82.19
LDFS w/ global 96.34 81.66 82.95
WaffleCLIP [30] 95.37 80.34 81.19
CLIP (ZS) 94.77 79.27 80.04
Analysis of Instance-conditional Loss. We compared our loss variant using averaged random descriptions, inspired by WaffleCLIP [30], denoted as $LDFS\_en$. However, we found no significant improvement over using a single description. Additionally, our local adaptation strategy substantially outperformed the global adaptation approach (i.e., LDFS w/ global). Our observations indicate that more semantically relevant and diverse descriptions lead to improved feature synthesis and performance gains. In contrast, simply adding or averaging stochastic directions does not yield comparable improvements. These findings underscore the importance of quality and diversity in descriptions rather than mere quantity or randomness.
@@
The Effect of Text Description. To explore if our method relies on users’ verbalization skills, we ablate different text descriptions in Table IX. It shows consistent efficacy despite variations in description format. It can also be seen that all prompts improve over the CoOp [45] baseline, with ”a photo of a class name with domain name background” achieving the best result. Moreover, Table X displays straightforward, uniform text descriptions used across all our datasets in our paper. The simplicity of these descriptions demonstrates that our method is easily applicable and user-friendly.
@@
Figure 11: Effect of number of shots on PACS and OfficeHome. The red and blue dash lines represent the performance of the model without using LDFS on PACS and OfficeHome.
Effect of Number of Shots. In Fig. 11, we plot the average performance of our LDFS framework using different few-shot settings (1, 4, 8, 10, and 16 shots per class) on the PACS [18] and OfficeHome [32] datasets. As can be seen, the performance of LDFS overall increases as the number of shots increases. All settings with LDFS outperform the CoOp baseline without LDFS, even when using just 1 shot per class.